# Breaking the Manual Annotation Bottleneck: Creating a Comprehensive Legal Case Criticality Dataset through Semi-Automated Labeling

### v1

@@ Abstract
Predicting case criticality helps legal professionals in the court system manage large volumes of case law. This paper introduces the Criticality Prediction dataset, a new resource for evaluating the potential influence of Swiss Federal Supreme Court decisions on future jurisprudence. Unlike existing approaches that rely on resource-intensive manual annotations, we semi-automatically derive labels leading to a much larger dataset than otherwise possible. Our dataset features a two-tier labeling system: (1) the LD-Label, which identifies cases published as Leading Decisions (LD), and (2) the Citation-Label, which ranks cases by their citation frequency and recency. This allows for a more nuanced evaluation of case importance. We evaluate several multilingual models, including fine-tuned variants and large language models, and find that fine-tuned models consistently outperform zero-shot baselines, demonstrating the need for task-specific adaptation. Our contributions include the introduction of this task and the release of a multilingual dataset to the research community.
Breaking the Manual Annotation Bottleneck: Creating a Comprehensive Legal Case Criticality Dataset through Semi-Automated Labeling
Ronja Stern 1thanks:    Equal contribution.                        Ken Kawamura 5
@@ 1 Introduction
Predicting the impact of legal cases is a critical task in the legal domain, as it aids professionals in the judicial system in navigating large volumes of case law. Despite its significance, the task of predicting the case criticality remains relatively under-explored. Existing approaches to evaluating the importance of legal cases are primarily manual, very resource-intensive and subject to the judgments of individual annotators. This paper introduces a novel dataset, licensed under CC BY 4.0, [^1] https://huggingface.co/datasets/rcds/swiss_criticality_prediction and a more challenging evaluation framework—Criticality Prediction—designed to predict the potential influence of Swiss Federal Court cases on future jurisprudence.
While prior work such as the Importance Prediction task proposed by Chalkidis et al. (2019) for European Court of Human Rights (ECtHR) cases focused on predicting importance on a defined scale using human-assigned labels, our approach employs algorithmically derived labels to evaluate case criticality. Our dataset introduces a two-tier labeling system: the LD-Label, a binary indicator of whether a case is published as a Leading Decision (LD), and the Citation-Label, a more nuanced categorization based on the frequency and recency-weighted citation counts of these decisions across subsequent cases. This automated and distinct formulation of "criticality" not only distinguishes critical cases from non-critical ones but also ranks them by their relative importance over time. The complexity of this dataset challenges even recent large language models (LLMs) such as GPT-3.5 (Brown et al., 2020).
Our contributions are threefold: (1) We propose a novel Criticality Prediction task that provides a more comprehensive and challenging evaluation of case law importance. (2) We release the datasets to the community, providing valuable resources for further research in legal NLP. (3) We evaluate several multilingual models of various sizes, including fine-tuned variants to set baselines.
@@ 2 Related Work
One of the most common text classification tasks in the legal domain is Legal Judgment Prediction (LJP), which involves predicting the outcome based on its facts. Researchers have leveraged diverse datasets with unique characteristics and annotations to analyze and predict case outcomes across various languages, jurisdictions, and input types (Feng et al., 2022; Aletras et al., 2016; Şulea et al., 2017; Medvedeva et al., 2018; Chalkidis et al., 2019; Niklaus et al., 2021, 2022; Semo et al., 2022). While LJP is focused on the outcome of individual cases, Importance Prediction shifts the focus toward assessing the broader significance of a case.
Chalkidis et al. (2019) introduced the Importance Prediction task using cases from the European Court of Human Rights (ECtHR). In this task, ECtHR provided scores that denote each case’s ‘importance’ to the common law. These scores, ranging from 1 (key case) to 4 (unimportant), were designed to help legal practitioners identify cases that play a crucial role in shaping jurisprudence. The labels reflect the long-term impact of a case on future rulings and the evolution of legal precedent. While the task is invaluable for identifying landmark cases, it relies on legal experts to assign the labels, making the process resource-intensive and potentially subject to subjective interpretations. Additionally, the ECtHR dataset is monolingual in English, whereas our Swiss dataset is multilingual. To our knowledge, no other study addresses a similar task to Criticality Prediction.
@@ 3 Task and Dataset
Task Name Train Labels Train Validation Labels Validation Test Labels Test
@@ C-1 C-2 C-3 C-4 C-1 C-2 C-3 C-4 C-1 C-2 C-3 C-4
3.1 Criticality
Understanding the legal framework of the Swiss Federal Supreme Court (SFSC) is key to defining criticality in Swiss case law. The SFSC shapes the legal landscape through its rulings, with a subset known as Leading Decisions published separately due to their influence on future decisions. We quantify case criticality using two labels: the binary LD-Label and the more granular Citation-Label.
The LD-Label is binary, categorizing cases as critical or non-critical. SFSC cases are labeled critical if also published as LD, reflecting their recognized importance within the Swiss legal system. We used regular expressions to extract SFSC cases from LD headers if present. SFSC cases not published as LD are labeled non-critical.
We developed the Citation-Label to provide a more granular measure. It counts how often each LD case is cited in SFSC cases, with older cases weighted less to prioritize recency. The score is calculated as: $score=count\times\frac{year-2002+1}{2023-2002+1}$. The count is the citation frequency, and the weighting reduces older cases’ influence. More details on the constants are in Appendix A. This score ranks LD cases, and we categorized them into four criticality levels—critical-1 (least critical) to critical-4—based on the 25th, 50th, and 75th percentiles.
Unlike prior approaches such as  Chalkidis et al. (2019), we explicitly incorporate temporal weighting to account for both the influence of a case and how its criticality shifts over time. Furthermore, our framework allows for the dynamic recalculation of scores and re-labeling of criticality as case law evolves. This adaptability ensures that our dataset reflects the ongoing changes in the legal system, whereas prior studies would require manual re-annotation as the law develops.
@@ 3.2 Criticality Prediction Task and Dataset
We see two distinct applications for these inputs in the Criticality Prediction task, illustrated in Figure 1. In the Case Prioritization task, only the facts are used as input. This produces a score indicating how critical or important a case is. The goal is to help prioritize cases, which could assist in determining which cases should be heard sooner or assigned to more experienced judges.
@@
Our dataset spans from 2002 to 2023, and was partitioned into training (2002-2015), validation (2016-2017), and test (2018-2022) sets, as outlined in Table 1. We allocated a relatively large test set to accommodate longitudinal studies, including the COVID-19 pandemic years. The dataset consists of 138,531 total cases, with 85,167 in German, 45,451 in French, and 7,913 in Italian. See Figure 2 for length distributions of facts and considerations and Appendix Table 4 for general dataset metadata.
@@ 4.1 Models
We evaluated the following models, various size variants: MiniLM (Wang et al., 2020), DistilmBERT (Sanh, 2019), mDeBERTa-v3 (He et al., 2021), XLM-R (Base and Large) (Conneau and Lample, 2019), X-MOD (Base) (Pfeiffer et al., 2022), SwissBERT (Vamvas et al., 2023), mT5 (Small and Base) (Xue et al., 2021), BLOOM (560M) (Scao et al., 2022), Legal-Swiss-RoBERTa and Legal Swiss Longformer (Base) (Rasiah et al., 2023), GPT-3.5 (Brown et al., 2020), and LLaMA-2 (Touvron et al., 2023).
We fine-tuned all models per task, using early stopping on the validation dataset. Due to resource constraints, further fine-tuning of GPT-3.5 and LLaMA-2 is reserved for future work, with their current performance serving as baseline results. SwissBERT and Legal-Swiss models were chosen for their Swiss-specific pretraining, while the other models were chosen for their multilingual capabilities, essential for our multilingual dataset.
We evaluated GPT-3.5 and LLaMA-2 0-shot following Chalkidis (2023), using one instruction and example as input. Samples were randomly selected from the validation set to prevent test set leakage for future evaluations especially for a closed model (GPT-3.5). To manage costs, we limited the validation set to 1000 samples. Our experiments focused solely on zero-shot classification due to the long input lengths. We show the prompts used in Appendix Figure 3 and Figure 4. We used the ChatCompletion API for GPT-3.5 (as of June 7, 2023), and ran LLaMA-2 locally with 4-bit quantization.
@@ 4.2 Metrics
We adopt the LEXTREME benchmark setup (Niklaus et al., 2023), and use hierarchical aggregation of macro-averaged F1 scores with the harmonic mean to emphasizes lower scores, promoting fairness across languages and input types. Scores are averaged over random seeds, languages (de, fr, it), and input types (facts or considerations), penalizing models with outlier low scores, encouraging consistent performance across all configurations.
@@ 5 Results
Table 3: Results using Macro F1, with highest values in bold. ’F’ and ’C’ refer to inputs from Facts or Considerations, while ’LD’ and ’C’ denote LD and Citation labels. Models marked with (*) are zero-shot LLMs evaluated on the validation set.
@@
We present results in Table 3, with standard deviations in Appendix Table 5 and scores on the validation dataset in Appendix Table 6. The best performance was achieved by XLM-RLarge, with an aggregate (Agg.) score of 37.1. SwissBERT also demonstrated competitive results, with an Agg. score of 34.8. Interestingly, larger models did not always outperform their smaller counterparts. For example, mT5Base and mT5Small both underperformed compared DistilmBERT in all configurations.
The Legal-Swiss models performed well in LD labels, particularly the Legal-Swiss-LFBase, which achieved the highest scores in LD-F and LD-C. However, their weak performance on Citation labels highlights the dataset’s complexity, even for domain-specific models.
LLMs such as GPT-3.5 and LLaMA-2 underperform fine-tuned models, underlining the need for specialized models for these tasks. The difference is largest in the LD labels where the small fine-tuned models always outperformed LLMs.
@@
Overall, there only seems to be a weak trend connecting higher percentage of a given language in the pretraining corpus leading to better downstream results in that language.
@@ 6 Conclusions and Future Work
This work introduced a novel Criticality Prediction task to assess the potential influence of Swiss Federal Court cases on future jurisprudence. Our approach utilizes algorithmically derived labels for a more comprehensive and challenging multilingual evaluation of case law importance compared to existing methods. We also released its multilingual dataset to advance research in legal NLP.
We conducted a comprehensive evaluation of our proposed Criticality Prediction task, comparing a diverse range of models, from smaller multilingual models to large language models like GPT-3.5. Our findings demonstrate that fine-tuned models consistently outperform their zero-shot LLMs, achieving superior macro-F1 scores. This underscores the importance of task-specific adaptation for optimizing performance in legal NLP applications.
Future studies could explore application of the Criticality Prediction task in other legal contexts by incorporating sources from different jurisdictions and languages. This would broaden the impact of this research, providing valuable insights into cross-jurisdictional case influences and enhancing the model’s adaptability to various legal systems.
@@
It is very difficult to estimate the importance of a case. By relying on proxies such as whether the case was converted to a leading decision (LD-label) and how often this leading decision was cited (Citation-label), we were able to create labels semi-automatically. While we discussed this with lawyers at length and implemented the solution we agreed on finally, this task remains somewhat artificial.
@@ Appendix C General Dataset Metadata
Table 4: Listing of cantons, courts, chambers, law-areas
@@ Appendix F Hyperparameters and Package Settings
We used a fixed learning rate of 1e-5 without tuning, running each experiment with three random seeds (1-3) and excluding seeds with high evaluation losses. Gradient accumulation was applied when GPU memory was insufficient to maintain a final batch size of 64. Training employed early stopping with a patience of 5 epochs, based on validation loss. To reduce costs, AMP mixed precision was used where it didn’t cause overflows (e.g., mDeBERTa-v3). The max-sequence length was set at 2048 for Facts and 4096 for Considerations.

### v2

# From Citations to Criticality: Predicting Legal Decision Influence in the Multilingual Swiss Jurisprudence

@@ Abstract
Many court systems are overwhelmed all over the world, leading to huge backlogs of pending cases. Effective triage systems, like those in emergency rooms, could ensure proper prioritization of open cases, optimizing time and resource allocation in the court system. In this work, we introduce the Criticality Prediction dataset, a novel resource for evaluating case prioritization. Our dataset features a two-tier labeling system: (1) the binary LD-Label, identifying cases published as Leading Decisions (LD), and (2) the more granular Citation-Label, ranking cases by their citation frequency and recency, allowing for a more nuanced evaluation. Unlike existing approaches that rely on resource-intensive manual annotations, we algorithmically derive labels leading to a much larger dataset than otherwise possible. We evaluate several multilingual models, including both smaller fine-tuned models and large language models in a zero-shot setting. Our results show that the fine-tuned models consistently outperform their larger counterparts, thanks to our large training set. Our results highlight that for highly domain-specific tasks like ours, large training sets are still valuable.
From Citations to Criticality:
Predicting Legal Decision Influence in the Multilingual Swiss Jurisprudence
Ronja Stern 1thanks:    Equal contribution.                        Ken Kawamura 5∗
@@ 1 Introduction
Predicting the impact of legal cases is a critical task in the legal domain, as it aids professionals in the judicial system prioritize and manage large volumes of case law more effectively, ensuring that critical cases receive the necessary attention. Despite its significance, the task of predicting the case criticality remains relatively under-explored. Existing approaches to evaluating the importance of legal cases are primarily manual, very resource-intensive and subject to the judgments of individual annotators. This paper introduces a novel dataset , licensed under CC BY 4.0, [^1] https://huggingface.co/datasets/rcds/swiss_criticality_prediction and a more challenging evaluation framework—Criticality Prediction—designed to predict the potential influence of Swiss Federal Court cases on future jurisprudence.
While prior work such as the Importance Prediction task for European Court of Human Rights (ECtHR) Chalkidis et al. (2019) cases focused on predicting importance on a defined scale using human-assigned labels, our approach employs algorithmically derived labels to evaluate case criticality. Our dataset introduces a two-tier labeling system: the LD-Label, a binary indicator of whether a case is published as a Leading Decision (LD), and the Citation-Label, a more nuanced categorization based on the frequency and recency-weighted citation counts of these decisions across subsequent cases. This distinct formulation of "criticality" not only distinguishes critical cases from non-critical ones but also ranks them by their relative importance over time. The complexity of this dataset challenges even recent LLMs such as ChatGPT.
Our contributions are threefold: (1) We propose a novel Criticality Prediction task that provides a more comprehensive and challenging evaluation of case law importance. (2) We release the datasets to the community, providing valuable resources for further research in legal NLP. (3) We evaluate several multilingual models of various sizes, including fine-tuned variants to establish baselines.
@@ 2 Related Work
One of the most common text classification tasks in the legal domain is Legal Judgment Prediction (LJP), which involves predicting outcome of a case based on its facts. Researchers have leveraged datasets with unique characteristics to analyze and predict case outcomes across various languages, jurisdictions, and input types (Feng et al., 2022; Aletras et al., 2016; Şulea et al., 2017; Medvedeva et al., 2018; Chalkidis et al., 2019; Niklaus et al., 2021, 2022; Semo et al., 2022). While LJP is focused on the outcome of individual cases, Importance Prediction shifts the focus toward assessing the broader significance of a case.
Chalkidis et al. (2019) introduced the Importance Prediction task using cases from the European Court of Human Rights (ECtHR). In this task, ECtHR provided scores that denote each case’s ‘importance’ to the common law. These scores, ranging from 1 (key case) to 4 (unimportant), were designed to help legal practitioners identify cases that play a crucial role in shaping jurisprudence. The labels reflect the long-term impact of a case on future rulings and the evolution of legal precedent. While the task is invaluable for identifying landmark cases, it relies on legal experts to assign the labels, making the process resource-intensive and potentially subject to subjective interpretations. Additionally, the ECtHR dataset contains only English samples, whereas our Swiss dataset is multilingual. To our knowledge, no other study has specifically addressed the task of Criticality Prediction.
@@ 3 Task and Dataset
Task Name Train Validation Test
Size Label Distribution Size Label Distribution Size Label Distribution
@@ C-1 C-2 C-3 C-4 C-1 C-2 C-3 C-4 C-1 C-2 C-3 C-4
3.1 Criticality
Understanding the legal framework of the Swiss Federal Supreme Court (SFSC) is key to defining criticality in Swiss case law. The SFSC shapes the legal landscape through its rulings, with a subset known as Leading Decisions published separately due to their influence on the interpretation of law. We quantify case criticality using two labels: the LD-Label and the Citation-Label.
The LD-Label is binary, categorizing cases as critical or non-critical. SFSC cases are labeled critical if also published as LD, reflecting their recognized importance within the Swiss legal system. We used regular expressions to extract SFSC cases published as LD. SFSC cases not published as LD are labeled non-critical.
We developed the Citation-Label for a more granular measure. It counts how often each LD case is cited in SFSC cases, with less weight on older cases to prioritize recency. The score is calculated as: $score=count\times\frac{year-2002+1}{2023-2002+1}$. The count is the citation frequency, and the weighting reduces older cases’ influence. This score ranks LD cases, and we categorized them into four criticality levels—critical-1 (least critical) to critical-4—based on the 25th, 50th, and 75th percentiles. More details on the constants are in Appendix A.
@@ 3.1 Criticality
Unlike a prior approach Chalkidis et al. (2019), which also used four categories, we explicitly incorporate temporal weighting to account for both the influence of a case and how its criticality shifts over time. Furthermore, our framework allows for the dynamic recalculation of scores and re-labeling of criticality as case law evolves. This adaptability ensures that our dataset reflects the ongoing changes in the legal system, whereas prior studies would require manual re-annotation as the law develops.
@@ 3.2 Criticality Prediction Task and Dataset
We see two distinct applications for these inputs in the Criticality Prediction task, illustrated in Figure 1. In the Case Prioritization task, only the facts are used as input. This produces a score indicating how critical or important a case is. The goal is to help prioritize cases, which could assist in determining which cases should be heard sooner or assigned to more experienced judges. Many court systems are overwhelmed all over the world leading to huge backlogs of pending cases. Effective triage systems, like those in emergency rooms, could ensure proper prioritization of open cases.
@@
Our dataset spans from 2002 to 2023, and is partitioned into train (2002-2015), dev (2016-2017), and test (2018-2022) sets, as outlined in Table 1. We allocated a relatively large test set to accommodate longitudinal studies, including the COVID-19 pandemic years. This setup ensures that the test set represents the most recent data for realistic evaluation Søgaard et al. (2021). The dataset consists of 138,531 total cases, with 85,167 in German, 45,451 in French, and 7,913 in Italian. See Figure 2 for length distributions of facts and considerations and Appendix Table 4 for general dataset metadata.
@@ 4.1 Models
We evaluated the following models: XLM-R (Base and Large) (Conneau and Lample, 2019), MiniLM (Wang et al., 2020), DistilmBERT (Sanh, 2019), mDeBERTa-v3 (He et al., 2021), X-MOD (Base) (Pfeiffer et al., 2022), SwissBERT (Vamvas et al., 2023), mT5 (Small and Base) (Xue et al., 2021), BLOOM (560M) (Scao et al., 2022), Legal-Swiss-RoBERTa and Legal Swiss Longformer (Base) (Rasiah et al., 2023), GPT-3.5 (Brown et al., 2020), and LLaMA-2 (Touvron et al., 2023).
We fine-tuned all models per task, using early stopping on the dev set. Due to resource constraints, further fine-tuning of GPT-3.5 and LLaMA-2 is reserved for future work, with their current performance serving as baseline results. SwissBERT and Legal-Swiss models were chosen for their Swiss-specific pretraining, while the other models were chosen for their multilingual capabilities, essential for our multilingual dataset.
We evaluated GPT-3.5 and LLaMA-2 in a zero-shot setting following Chalkidis (2023). Samples were randomly selected from the validation set to prevent test set leakage for future evaluations especially for a closed model (GPT-3.5). To manage costs, we limited the validation set to 1000 samples. Our experiments focused on zero-shot classification due to the long input lengths. The prediction labels were determined using regular expressions to match model outputs to class labels. We show the prompts used in Appendix Figure 4 and Figure 5.
@@ 4.2 Metrics
We adopt the LEXTREME benchmark setup (Niklaus et al., 2023), and use aggregation of macro-averaged F1 scores with the harmonic mean to emphasize lower scores, promoting fairness across languages and input types. Scores are averaged over random seeds, languages (de, fr, it), and input types (facts or considerations), encouraging consistent performance across configurations.
@@ 5 Results
Table 3: Results using Macro F1, with highest values in bold. ’LD’ and ’C’ denote LD and Citation labels, while ’F’ and ’C’ refer to inputs from Facts or Considerations. Models marked with (*) are zero-shot LLMs evaluated on the validation set.
@@
We present results in Table 3, with standard deviations in Appendix Table 5 and scores on the validation dataset in Appendix Table 6. The best performance was achieved by XLM-RLarge, with an aggregate (Agg.) score of 37.1. SwissBERT also demonstrated competitive results, with an Agg. score of 34.8. Interestingly, larger models did not always outperform their smaller counterparts. For example, mT5Base and mT5Small both underperformed DistilmBERT in all configurations. The Legal-Swiss models performed well in LD labels, particularly the Legal-Swiss-LFBase, which achieved the highest scores in LD-F and LD-C. However, their weak performance on Citation labels highlights the dataset’s complexity, even for domain-specific models. GPT-3.5 and LLaMA-2 underperform fine-tuned models, underlining the need for specialized models for these tasks. The difference is largest in the LD labels where the small fine-tuned models always outperformed LLMs despite their huge parameter number difference (e.g., MiniLM with 118M parameters versus GPT-3.5 with 175B).
@@
Overall, the connection between the proportion of a language in the pretraining corpus and the model’s downstream performance in that language appears weak. While some models showed stronger results in languages with higher representation in their pretraining data, this trend was inconsistent across all models and languages.
@@ 6 Conclusions and Future Work
This work introduced a novel Criticality Prediction task to assess the potential influence of SFSC cases on future jurisprudence. Our approach utilizes leading decision citation for a more comprehensive and challenging multilingual evaluation of case law importance compared to existing methods. We also released its multilingual dataset for a community.
We conducted a comprehensive evaluation of our Criticality Prediction task, comparing a range of models, from smaller multilingual models to LLMs like GPT-3.5. Our findings show that small fine-tuned models consistently outperform zero-shot LLMs, highlighting the importance of task-specific adaptation in legal NLP applications.
Future studies could apply the Criticality Prediction task in other legal contexts by incorporating sources from different jurisdictions and languages. This approach would broaden the research’s impact and enhance the model’s adaptability across different legal systems.
@@
It is very difficult to estimate the importance of a case. By relying on proxies such as whether the case was converted to a leading decision (LD-label) and how often this leading decision was cited (Citation-label), we were able to create labels algorithmically. While we discussed this with lawyers at length and implemented the solution we agreed on finally, this task remains somewhat artificial. Additionally, it is worth noting that LLMs can be sensitive to prompt formats and the order of answer options, leading to inconsistent outputs as highlighted in recent research Webson and Pavlick (2022); Elazar et al. (2021). Incorporating varied prompt designs into future studies could strengthen our approach further.
@@
Acknowledgments
We thank Luca Rolshoven and Mara Häusler for their thoughtful feedback on the manuscript, and the anonymous reviewers for their insightful and constructive comments.
@@ Appendix A Weighting Formula for Citation-Label
We did not use more than four bins for the Citation-Label to ensure enough examples also for the smallest class while still adding more granularity compared to the LD-Label.
@@ Appendix B Dataset Example
Figure 3: Example of an input, target and corresponding metadata
Figure 3 shows an example input, target and corresponding metadata.
@@ Appendix C General Dataset Metadata
Table 4 shows metadata for cantons, courts, chambers law-areas and languages.
Table 4: Metadata for cantons, courts, chambers, law-areas, and languages
@@ Appendix E Zero-shot Prompts
Figures 4 and 5 show the zero-shot prompts we used for the Case Prioritization and Leading Decision Selection tasks respectively.
@@ Appendix F Hyperparameters and Package Settings
We used a fixed learning rate of 1e-5 without tuning, running each experiment with three random seeds (1-3) and excluding seeds with high evaluation losses. Gradient accumulation was applied when GPU memory was insufficient to maintain a final batch size of 64. Training employed early stopping with a patience of 5 epochs, based on validation loss. To reduce costs, AMP mixed precision was used where it didn’t cause overflows (e.g., mDeBERTa-v3). The max-sequence length was set at 2048 for Facts and 4096 for Considerations. For LLM evaluations, we used the ChatCompletion API for GPT-3.5 (as of June 7, 2023), and ran LLaMA-2 locally with 4-bit quantization.
@@ Appendix H Additional Results
Table 5 shows results with standard deviations on the test set. Table 6 shows scores on the validation set.